# Continual Learning of Large Language Models: A Comprehensive Survey

### v2

@@ Abstract.
The challenge of effectively and efficiently adapting statically pre-trained Large Language Models (LLMs) to ever-evolving data distributions remains predominant. When tailored for specific needs, pre-trained LLMs often experience significant performance degradation in previous knowledge domains – a phenomenon known as “catastrophic forgetting”. While extensively studied in the continual learning (CL) community, this problem presents new manifestations in the realm of LLMs. In this survey, we provide a comprehensive overview and detailed discussion of the current research progress on LLMs within the context of CL. Besides the introduction of the preliminary knowledge, this survey is structured into four main sections: we first describe an overview of continually learning LLMs, consisting of two directions of continuity: vertical continuity (or vertical continual learning), i.e., continual adaptation from general to specific capabilities, and horizontal continuity (or horizontal continual learning), i.e., continual adaptation across time and domains (Section 3). Following vertical continuity, we summarize three stages of learning LLMs in the context of modern CL: Continual Pre-Training (CPT), Domain-Adaptive Pre-training (DAP), and Continual Fine-Tuning (CFT) (Section 4). Then we provide an overview of evaluation protocols for continual learning with LLMs, along with the current available data sources (Section 5). Finally, we discuss intriguing questions pertaining to continual learning for LLMs (Section 6). This survey sheds light on the relatively understudied domain of continually pre-training, adapting, and fine-tuning large language models, suggesting the necessity for greater attention from the community. Key areas requiring immediate focus include the development of practical and accessible evaluation benchmarks, along with methodologies specifically designed to counter forgetting and enable knowledge transfer within the evolving landscape of LLM learning paradigms. The full list of papers examined in this survey is available at https://github.com/Wang-ML-Lab/llm-continual-learning-survey.
@@
copyright: nonejournal: CSURccs: Computing methodologies Lifelong machine learningccs: Computing methodologies Natural language processingccs: Computing methodologies Neural networks
@@ 1. Introduction
Recent advances in large language models (LLMs) have demonstrated considerable potential for achieving artificial general intelligence (AGI) (radford2019language, ; brown2020language, ; achiam2022chatgpt, ; achiam2023gpt, ; chowdhery2023palm, ; anil2023palm, ; touvron2023llama, ; touvron2023llama2, ; jiang2024empowering, ). Researchers have observed that complex abilities such as multi-step reasoning, few-shot in-context learning, and instruction following improve as the scale of parameter size increases (wei2022chain, ; wei2022emergent, ; yao2024tree, ; wei2021finetuned, ; min2022rethinking, ). The development of LLMs is impactful and revolutionary, prompting machine learning practitioners to reconsider traditional computational paradigms for once-challenging human-level tasks such as question answering, machine translation, and dialogue systems (kwok2001scaling, ; bahdanau2014neural, ; deng2023survey, ). However, LLMs are typically trained on static, pre-collected datasets encompassing general domains, leading to gradual performance degradation over time (loureiro2022timelms, ; jang2022towards, ; jin2022lifelong, ; jang2022temporalwiki, ; amba2021dynamic, ; dhingra2022time, ) and across different content domains (gupta2023continual, ; jin2022lifelong, ; ke2022continual-train, ; sun2020ernie, ; cossu2022continual, ; gururangan2022demix, ; qin2023recyclable, ; chen2023lifelong, ; qin2022elle, ). Additionally, a single pre-trained large model cannot meet every user need and requires further fine-tuning (weyssow2023usage, ; winata2023overcoming, ; zheng2023learn, ; winata2023overcoming, ; biderman2023pythia, ; zheng2023learn, ; bai2023enhancing, ; ke2021achieve, ; wei2022circle, ; qin2021lfpt5, ; chen2024parameterizing, ). While one potential solution is re-collecting pre-training data and re-training models with additional specific needs, this approach is prohibitively expensive and impractical in real-world scenarios.
@@
To efficiently adapt LLMs to downstream tasks while minimizing performance degradation on previous knowledge domains, researchers employ the methodology of continual learning, also known as lifelong learning or incremental learning (pentina2016theoretical, ; chen2018lifelong, ; van2022three, ; wang2024comprehensive, ). Continual learning, inspired by the incremental learning pattern observed in human brains (mcclelland1995there, ; kandel2000principles, ; pallier2003brain, ; yang2009stably, ; constantinescu2016organizing, ; olafsdottir2018role, ; liu2019human, ; mccaffary2021towards, ), involves training machine learning models sequentially on a series of tasks with the expectation of maintaining performance across all tasks (kirkpatrick2017overcoming, ; li2017learning, ; zenke2017continual, ; riemer2018learning, ; buzzega2020dark, ; garg2023in, ; ebrahimi2020adversarial, ; ebrahimi2019uncertainty, ). Throughout training, models have limited or no access to previous data, posing a challenge in retaining past knowledge as optimization constraints from unseen previous data are absent during current-task learning (li2017learning, ; smith2023closer, ; hayes2020lifelong, ; lomonaco2020rehearsalfree, ; chaudhry2019tiny, ; riemer2018learning, ; buzzega2020dark, ; shi2024unified, ). This challenge, known as catastrophic forgetting (mccloskey1989catastrophic, ), has been a central focus in continual learning research since its inception. Over the years, researchers have explored various techniques to mitigate forgetting in machine learning models. These include replay-based methods (chaudhry2019tiny, ; schwarz2018progress, ; riemer2018learning, ; buzzega2020dark, ; shi2024unified, ), parameter regularization (kirkpatrick2017overcoming, ; ritter2018online, ; aljundi2018memory, ; sprechmann2018memory, ), and model architecture expansion (ramesh2021model, ; wang2022coscl, ). Together, these techniques have significantly advanced the goal of achieving zero forgetting in continual learning across diverse tasks, model architectures, and learning paradigms.
@@
In the context of training and adapting LLMs sequentially, the significance of CL is undergoing semantic shifts of its own as well. To better highlight this ongoing shift, in this survey, we provide a comprehensive overview and detailed discussion of the current research progress on LLMs within the context of CL. For the general picture of continually learning LLMs, we for the first time divide it into two directions of continuity that need to be addressed by practitioners (Section 3):
@@
The explicit separation of vertical and horizontal CL extends beyond a trivial modification of existing CL types, such as domain-incremental learning, which might be considered analogous to horizontal continuity. It offers a robust conceptual framework for analyzing and describing complex learning paradigms in continual LLMs. For example, Recyclable Tuning aims to preserve both vertical and horizontal continuity simultaneously (qin2023recyclable, ), and future designs could include zigzag CL, alternating between horizontal and vertical CL.
@@
In Fig. 1, following vertical continuity, we delineate three key stages of LLM learning within modern CL: Continual Pre-Training (CPT), Domain-Adaptive Pre-training (DAP), and Continual Fine-Tuning (CFT) (Section 4). In CPT, existing research primarily investigates three types of distributional shifts: temporal, content-level, and language-level. Each presents distinct focuses and challenges. In DAP, while it is primarily seen as the procedure of preparing LLMs for downstream tasks, CL evaluation and techniques are frequently utilized. However, there is a noticeable lack of diversity in these techniques, considering the maturity of the conventional CL community. In CFT, our focus is on the emerging field of learning LLMs, covering topics such as Continual Instruction Tuning (CIT), Continual Model Refinement (CMR), Continual Model Alignment (CMA), and Continual Multimodal LLMs (CMLLMs). Next, we present a compilation of publicly available evaluation protocols and benchmarks (Section 5). We conclude our survey with a discussion covering emergent properties of continual LLMs, changes in the roles of conventional CL types and memory constraints within the context of continual LLMs, and prospective research directions for this subject (Section 6).
@@
In summary, this paper provides a comprehensive view of existing continual learning studies for LLMs in detail, which significantly distinguishes itself from existing literature on related topics (biesialska2020continual, ; ke2023continual, ; wang2024comprehensive, ; wu2024continual, ; yang2024recent, ). Our survey highlights the underexplored research area of continually developing LLMs, especially in the field of continual pre-training (CPT) and domain adaptive pre-training (DAP). We emphasize the needs for increased attention from the community, with urgent needs including the development of practical, accessible, and widely acknowledged evaluation benchmarks. Additionally, methodologies need to be tailored to address forgetting in emerging large language model learning paradigms. We hope this survey can provide a systematic and novel view of continual learning in the rapidly-changing field of LLMs and can help the continual learning community contribute to the challenging goals of developing LLMs in a more efficient, reliable, and sustainable manner (jang2022temporalwiki, ; su2023efficient, ; xie2023efficient, ; Cao2023InstructMol, ; attanasio2023worth, ).
Organization. The rest of this paper is organized as follows. We will first start by introducing the background and preliminaries of large language models and continual learning in Section 2. Then we present the overview of continual learning in the modern era of large language models in Section 3. Vertically, it can be roughly divided into three stages of continual training LLMs, and we will present a one-by-one survey of each stage in Section 4. In Section 4.3, the unique aspects of continual fine-tuning LLMs will be introduced, including continual instruction tuning (Section 4.3.3), continual model refinement (Section 4.3.4), continual model alignment (Section 4.3.5), and continual multimodal large language models (Section 4.3.6). In Section 5, we give an inclusive introduction to the evaluation protocols and benchmarks of continual learning for LLMs that are publicly available. Finally, in Section 6, we present a series of discussion of the role of continual learning in the era of large language models, including emergent abilities in large-scale continual LLMs (Section 6.1), three types of continual learning (Section 6.2), roles of memory in continual learning of LLMs (Section 6.3), and prospective future directions (Section 6.4).
2. Preliminaries
In this section, we provide an overview of the fundamental concepts of large language models (LLMs) and continual learning (CL) We begin by introducing the notation used in this paper. Subsequently, we discuss the pre-training and downstream adaptation of LLMs, as well as mainstream LLM families (Section 2.1), followed by an introduction to basic continual learning techniques studied by the community (Section 2.2).
@@ 2.1.2. Adaptation of LLMs
After pre-training, LLMs need to be effectively adapted to better serve downstream tasks. A series of adaptation methods have been proposed for specific objectives. Due to the fact that LLMs primarily focus on generating linguistically coherent text during pre-training, their performance may not necessarily align with the actual needs of human users or conform to human values, preferences, and principles. Additionally, due to issues such as the timeliness of pre-training data, LLMs may also encounter knowledge cutoff or fallacy issues. Therefore, instruction tuning, model refinement, and model alignment have been proposed to address these issues (zhang2024instruction, ; ouyang2022rlhf, ; rafailov2024dpo, ; de2021editing, ). Below are the formal definitions of the three adaptation tasks for LLMs.
Definition 2.0 (Instruction Tuning, IT).
Remark 0.
Definition 2.0 (Model Refinement, MR).
@@ 2.2. Continual Learning
Humans gradually accumulate knowledge and skills across tasks without significant performance decline on previous tasks (mcclelland1995there, ; kandel2000principles, ; pallier2003brain, ; yang2009stably, ; constantinescu2016organizing, ; olafsdottir2018role, ; liu2019human, ; mccaffary2021towards, ). In contrast, machine learning models are usually data-centric, minimizing the training loss on the subsequent tasks will cause the model fail on the old ones, which phenomenon is phrased as “catastrophic forgetting”. Addressing this challenge is a focal point in continual learning research. The problem of efficiently adapting models to a sequence of tasks without forgetting is extensively studied in the continual learning community (pentina2016theoretical, ; chen2018lifelong, ; van2022three, ; wang2024comprehensive, ). These studies are typically conducted under the following memory constraint of CL.
Definition 2.0 (Memory Constraint of Continual Learning).
Remark 0.
In early stages of CL, works mostly focused on the strong memory constraint (kirkpatrick2017overcoming, ; li2017learning, ; aljundi2018memory, ; lomonaco2020rehearsalfree, ); as the research field progresses, more focus was put on relaxing the memory constraint to a small buffer for replay (rebuffi2017icarl, ; chaudhry2019tiny, ; buzzega2020dark, ; shi2024unified, ); some modern CL works completely discard the memory constraint but put focus on the computational budget (cai2021online, ; prabhu2023online, ; verwimp2024continual, ).
2.2.1. Three Types of Continual Learning
Definition 2.0 (Task-Incremental Learning, TIL).
Suppose $T$ task distributions $\{{\mathcal{T}}_{t}\}_{t=1}^{T}$ come in as a sequence, where ${\mathcal{T}}_{t}$ denotes the joint distribution over the $t$-th task’s input space and the label space $({\mathcal{X}}_{t},{\mathcal{Y}}_{t})$. Denote ${\mathcal{X}}\triangleq\bigcup_{t=1}^{T}{\mathcal{X}}_{t}$ and ${\mathcal{Y}}\triangleq\bigcup_{t=1}^{T}{\mathcal{Y}}_{t}$ as the union of the input and label spaces, respectively. Under the memory constraint defined in Definition 2.3, Task-Incremental Learning (TIL) aims to find the optimal hypothesis $h^{*}:{\mathcal{X}}\times[T]\rightarrow{\mathcal{Y}}$ that satisfies:
Definition 2.0 (Domain-Incremental Learning, DIL).
Suppose $T$ domain distributions $\{{\mathcal{D}}_{t}\}_{t=1}^{T}$ come in as a sequence, where ${\mathcal{D}}_{t}$ denotes the $t$-th joint distribution over the shared input space and label space $({\mathcal{X}},{\mathcal{Y}})$. Under the memory constraint defined in Definition 2.3, Domain-Incremental Learning (DIL) aims to find the optimal hypothesis $h^{*}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ that satisfies:
Definition 2.0 (Class-Incremental Learning, CIL).
Suppose $T$ task distributions $\{{\mathcal{T}}_{t}\}_{t=1}^{T}$ come in as a sequence, where ${\mathcal{T}}_{t}$ denotes the joint distribution over the $t$-th task’s input space and the label space $({\mathcal{X}}_{t},{\mathcal{Y}}_{t})$. Denote ${\mathcal{X}}\triangleq\bigcup_{t=1}^{T}{\mathcal{X}}_{t}$ and ${\mathcal{Y}}\triangleq\bigcup_{t=1}^{T}{\mathcal{Y}}_{t}$ as the union of the input and label spaces, respectively. Under the memory constraint defined in Definition 2.3, Class-Incremental Learning (CIL) aims to find the optimal hypothesis $h^{*}:{\mathcal{X}}\rightarrow[T]\times{\mathcal{Y}}$ that satisfies:
Remark 0.

### v3

@@ Abstract.
The challenge of effectively and efficiently adapting statically pre-trained Large Language Models (LLMs) to ever-evolving data distributions remains predominant. When tailored for specific needs, pre-trained LLMs often suffer from significant performance degradation in previous knowledge domains – a phenomenon known as “catastrophic forgetting”. While extensively studied in the Continual Learning (CL) community, this problem presents new challenges in the context of LLMs. In this survey, we provide a comprehensive overview and detailed discussion of the current research progress on LLMs within the context of CL. Besides the introduction of the preliminary knowledge, this survey is structured into four main sections: we first describe an overview of continually learning LLMs, consisting of two directions of continuity: vertical continuity (or vertical continual learning), i.e., continual adaptation from general to specific capabilities, and horizontal continuity (or horizontal continual learning), i.e., continual adaptation across time and domains (Section 3). Following vertical continuity, we summarize three stages of learning LLMs in the context of modern CL: Continual Pre-Training (CPT), Domain-Adaptive Pre-training (DAP), and Continual Fine-Tuning (CFT) (Section 4). We then provide an overview of evaluation protocols for continual learning with LLMs, along with currently available data sources (Section 5). Finally, we discuss intriguing questions related to continual learning for LLMs (Section 6). This survey sheds light on the relatively understudied domain of continually pre-training, adapting, and fine-tuning large language models, suggesting the necessity for greater attention from the community. Key areas requiring immediate focus include the development of practical and accessible evaluation benchmarks, along with methodologies specifically designed to counter forgetting and enable knowledge transfer within the evolving landscape of LLM learning paradigms. The full list of papers examined in this survey is available at https://github.com/Wang-ML-Lab/llm-continual-learning-survey.
@@
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: CSURccs: Computing methodologies Lifelong machine learningccs: Computing methodologies Natural language processingccs: Computing methodologies Neural networks
@@ 1. Introduction
Recent advances in large language models (LLMs) have demonstrated considerable potential for achieving artificial general intelligence (AGI) (radford2019language, ; brown2020language, ; achiam2022chatgpt, ; achiam2023gpt, ; chowdhery2023palm, ; anil2023palm, ; touvron2023llama, ; touvron2023llama2, ). Researchers have observed that complex abilities such as multi-step reasoning, few-shot in-context learning, and instruction following improve as the scale of parameter size increases (wei2022chain, ; wei2022emergent, ; yao2024tree, ; wei2021finetuned, ; min2022rethinking, ). The development of LLMs is impactful and revolutionary, prompting machine learning practitioners to reconsider traditional computational paradigms for once-challenging human-level tasks. However, LLMs are typically trained on static, pre-collected datasets encompassing general domains, leading to gradual performance degradation over time (loureiro2022timelms, ; jang2022towards, ; jin2022lifelong, ; jang2022temporalwiki, ; amba2021dynamic, ; dhingra2022time, ) and across different content domains (gupta2023continual, ; jin2022lifelong, ; ke2022continual-train, ; sun2020ernie, ; cossu2022continual, ; gururangan2022demix, ; qin2023recyclable, ; chen2023lifelong, ; qin2022elle, ). Additionally, a single pre-trained large model cannot meet every user need and requires further fine-tuning (weyssow2023usage, ; winata2023overcoming, ; zheng2023learn, ; winata2023overcoming, ; biderman2023pythia, ; zheng2023learn, ; bai2023enhancing, ; ke2021achieve, ; wei2022circle, ; qin2021lfpt5, ; chen2024parameterizing, ). While one potential solution is re-collecting pre-training data and re-training models with additional specific needs, this approach is prohibitively expensive and impractical in real-world scenarios.
@@
To efficiently adapt LLMs to downstream tasks while minimizing performance degradation on previous knowledge domains, researchers employ the methodology of Continual Learning (CL), also known as lifelong learning or incremental learning (pentina2016theoretical, ; chen2018lifelong, ; van2022three, ; wang2024comprehensive, ). Inspired by the incremental learning pattern observed in human brains (mcclelland1995there, ; kandel2000principles, ; pallier2003brain, ; mccaffary2021towards, ), CL trains machine learning models sequentially on a series of tasks with the expectation of maintaining performance across all tasks (kirkpatrick2017overcoming, ; li2017learning, ; ebrahimi2020adversarial, ; ebrahimi2019uncertainty, ). Throughout training, models have limited or no access to previous data, posing a challenge in retaining past knowledge as optimization constraints from unseen previous data are absent during current-task learning (li2017learning, ; lomonaco2020rehearsalfree, ; shi2024unified, ). This challenge, known as catastrophic forgetting (mccloskey1989catastrophic, ), has been a central focus in continual learning research since its inception. Over the years, researchers have explored various techniques to mitigate forgetting. These include replay-based methods (chaudhry2019tiny, ; schwarz2018progress, ; shi2024unified, ), parameter regularization (kirkpatrick2017overcoming, ; ritter2018online, ; aljundi2018memory, ), and model architecture expansion (ramesh2021model, ; wang2022coscl, ). Together, these techniques have significantly advanced the goal of achieving zero forgetting in continual learning across diverse tasks, model architectures, and learning paradigms.
@@
In the context of training and adapting LLMs sequentially, the significance of CL is undergoing semantic shifts of its own as well. To highlight this ongoing shift, in this paper, we provide a comprehensive overview and detailed discussion of the current research progress on continual LLMs. For the general picture of continual LLMs, we for the first time divide it into two directions of continuity that need to be addressed by practitioners (details in Section 3):
@@
Importantly, separating vertical and horizontal CL transcends mere modification of existing paradigms, like domain-incremental learning, which aligns with horizontal continuity. This distinction offers a robust framework for analyzing complex CL paradigms in language models. For instance, Recyclable Tuning preserves both vertical and horizontal continuity simultaneously (qin2023recyclable, ), and future designs might include zigzagging between horizontal and vertical CL.
@@
In Fig. 1, following vertical continuity, we delineate three key stages of LLM learning within modern CL: Continual Pre-Training (CPT), Domain-Adaptive Pre-training (DAP), and Continual Fine-Tuning (CFT) (details in Section 4). In CPT, existing research primarily investigates three types of distributional shifts: temporal, content-level, and language-level. Each presents distinct focuses and challenges. In DAP, CL evaluation and techniques are frequently utilized. However, there is a noticeable lack of diversity in these techniques, considering the maturity of the conventional CL community. In CFT, our focus is on the emerging field of learning LLMs, covering topics such as Continual Instruction Tuning (CIT), Continual Model Refinement (CMR), Continual Model Alignment (CMA), and Continual Multimodal LLMs (CMLLMs). Next, we present a compilation of publicly available evaluation protocols and benchmarks (details in Section 5). We conclude our survey with a discussion covering emergent properties of continual LLMs, changes in the roles of conventional CL types and memory constraints within the context of continual LLMs, and prospective research directions for this subject (details in Section 6).
@@
In summary, this survey provides a comprehensive review of existing continual learning studies for LLMs, which significantly distinguishes itself from existing literature on related topics (biesialska2020continual, ; ke2023continual, ; wang2024comprehensive, ; wu2024continual, ; yang2024recent, ). Our survey highlights the underexplored research area of continually developing LLMs, especially in the field of CPT and DAP. We emphasize the needs for increased attention from the community, including the development of practical, accessible, and widely acknowledged evaluation benchmarks. Additionally, methodologies need to be tailored to address forgetting in emerging LLM learning paradigms. We hope this survey can provide a systematic and novel perspective of continual learning in the rapidly-changing field of LLMs and can help the continual learning community contribute to the challenging goals of developing LLMs in a more efficient, reliable, and sustainable manner (jang2022temporalwiki, ; su2023efficient, ; xie2023efficient, ; Cao2023InstructMol, ; attanasio2023worth, ).
2. Background and Related Work
@@ 2.1. Large Language Models
Primarily built on the transformer architecture, pre-trained language models (PLMs) have established a universal hidden embedding space through extensive pre-training on large-scale unlabeled text corpora (devlin2018bert, ; liu2019roberta, ; raffel2020exploring, ). By scaling parameters to billions or even hundreds of billions and training on massive text datasets (kaplan2020scaling, ; hoffmann2022training, ), PLMs not only demonstrate superior language understanding and generation capabilities but also manifest emergent abilities such as in-context learning, instruction following, and multi-step reasoning (wei2022chain, ; wei2022emergent, ; yao2024tree, ; wei2021finetuned, ; min2022rethinking, ). These larger models are commonly referred to as Large Language Models (LLMs). For more detailed introduction, please refer to Appendix A.1.
@@ 2.1.1. Pre-training of LLMs
There are two popular pre-training paradigms for LLMs. (1) Decoder-only models typically employ auto-regressive language modeling (LM) tasks during pre-training, including the GPT family (radford2019language, ; brown2020language, ; achiam2022chatgpt, ; achiam2023gpt, ), Gemini family (team2023gemini, ; reid2024gemini, ), and the open-source Llama family (touvron2023llama, ; touvron2023llama2, ). Specifically, given a sequence of tokens ${\bm{x}}=[x_{1},x_{2},\cdots,x_{N}]$, LM predicts the next token $x_{t}$ autoregressively based on all preceding tokens ${\bm{x}}_{<t}=[x_{1},x_{2},\cdots,x_{t-1}]$, and trains the entire network by minimizing the negative log-likelihood $-\sum^{N}_{t=1}\log P(x_{t}|{\bm{x}}_{<t})$, where $P(x_{1}|{\bm{x}}_{<1})\triangleq P(x_{1})$ is the unconditional probability estimation of the first token. (2) Encoder-only models, e.g., BERT (devlin2018bert, ; liu2019roberta, ), use masked language modeling (MLM) as a common pre-training objective. In MLM, for the input sequence ${\bm{x}}$, a subset of input tokens $m({\bm{x}})$ are masked and replaced with the special [MASK] token. The pre-training goal is to utilize the unmasked parts ${\bm{x}}_{\backslash m({\bm{x}})}$ to predict the masked portions $m({\bm{x}})$. In summary, the overarching goal of MLM is to minimize the negative log-likelihood $-\sum_{\widehat{x}\in m({\bm{x}})}{\rm log}\,P(\widehat{x}|{\bm{x}}_{\backslash m({\bm{x}})})$.
@@ 2.1.2. Adaptation of LLMs
LLMs are primarily trained to generate linguistically coherent text. However, this training may not align with human values, preferences, or practical needs. Furthermore, the pre-training data can be outdated, leading to knowledge cutoffs or inaccuracies. To address these issues, various computational paradigms such as Instruction Tuning (IT) (zhang2024instruction, ), Model Refinement (MR) (de2021editing, ), and Model Alignment (MA) (ouyang2022rlhf, ; rafailov2024dpo, ) have been proposed. These approaches adapt LLMs to better meet diverse downstream tasks and user requirements.
Numerous studies show that Instruction Tuning (IT) can notably improve LLMs’ ability to follow textual instructions (zhang2024instruction, ; wei2021finetuned, ; jiang2024instructiontuned, ; sanh2022multitask, ; ouyang2022rlhf, ), leveraging the pre-existing knowledge within LLMs to bridge the gap between general and task-specific performance (wei2022finetuned, ). Recent works like WizardLM (xu2023wizardlm, ) and CodecLM (wang2024codeclm, ) further tailor synthetic data to steer LLMs’ behavior through IT. Additionally, IT enhances the interaction between humans and LLMs, providing a more natural interface and aligning LLM outputs more closely with human expectations and preferences (luo2023empirical, ). LLMs make mistakes, such as inaccurate translations or outdated information (de2021editing, ). Directly fine-tuning the model to correct these mistakes may disrupt its performance on previously learned tasks. To overcome these challenges, Model Refinement (MR) is proposed to rectify the model’s errors while preserving its performance on other inputs, with only moderate computing resources (sinitsin2020editable, ; de2021editing, ; fast_edit, ; hase2021language, ; huang2023transformer, ; mitchell2022memory, ; hartvigsen2023aging, ). Model Alignment (MA) ensures AI systems’ actions and outputs align with human values, ethics, and preferences (ouyang2022rlhf, ; rafailov2024dpo, ). MA can be broadly categorized into two types: Reinforcement Learning-based (RL-based) and Supervised Learning-based (SL-based). RL-based approaches (ouyang2022rlhf, ; schulman2017proximal, ) are trained to make decisions reinforced by human feedback, using a reward system to guide them towards desirable outcomes. In contrast, SL-based approaches (hendrycks2023aligning, ; rafailov2024dpo, ; ji2024ai, ) directly train models on datasets of human preferences, aligning their output with demonstrated human values.
@@ 2.2. Continual Learning
Humans can accumulate knowledge and skills across tasks without significant performance decline on previous tasks (mcclelland1995there, ; kandel2000principles, ; pallier2003brain, ; mccaffary2021towards, ). In contrast, machine learning models, which are typically data-centric, often experience performance degradation on old tasks when trained on new ones, a phenomenon known as “catastrophic forgetting.” The challenge of adapting models to a sequence of tasks without forgetting, especially when little to no past data can be preserved, is extensively studied in the continual learning community (pentina2016theoretical, ; chen2018lifelong, ; van2022three, ; wang2024comprehensive, ). For formal definitions, a detailed introduction to the three CL scenarios and techniques, please refer to Appendix A.2.
2.2.1. Types of Continual Learning
To lay the groundwork for subsequent discussions (as illustrated in Table 3 and Section 6.2), we follow the conceptual framework proposed by (van2022three, ; kim2022theoretical, ; wang2024comprehensive, ). There are three primary types of continual learning scenarios: (i) Task-Incremental Learning (TIL), where task indices are available to the model during inference (li2017learning, ; kirkpatrick2017overcoming, ); (ii) Domain-Incremental Learning (DIL), where the model learns a sequence of tasks with the same formulation but without task indices during inference (shi2024unified, ); and (iii) Class-Incremental Learning (CIL), where the model learns new classes of data during training (rebuffi2017icarl, ; kim2022theoretical, ).